OpenStudy (anonymous):

In a class of 160 students, 90 are taking math, 78 are taking science, and 62 are taking both math and science. What is the probability of randomly choosing a student who is not taking science?

OpenStudy (anonymous):

is 17.5% an option?

OpenStudy (anonymous):

Number of student taking only math = 90 - 62 = 28 Number of student taking only science = 78 - 62 = 16 Number of student taking math and science = 62 Number of student not taking science = 160 - 16 - 62 = 82 Probability of choosing student not taking science = 82 / 160